Assume that as a consequence of government directives to shut down in-house dining at restaurants, all but one restaurant in an area closes (but, to keep this problem from becoming awful, continues to operate in a competitive product market as if by magic). Given the following labor supply and demand equations, find the quantity of labor that will be employed by the restaurant and the wage they will pay.

w=1+0.25L

w=20-L

(Note that the supply equation implies MWC = 1 + 0.5L.)
Answer #1

Monopsony eqm,

Only one buyer, with large number of suppliers

Demand curve cuts MFC curve

MFC : Marginal factor cost

So, 20-L= 1+.5L

19 = 1.5L

L* = 19/1.5 = 12.667

finding Wages from supply Curve

w* = 1+.25*12.667

= 4.1667